OpenStudy (snickers1998):

What is the range of g(x)=3(2)^x-1

OpenStudy (mathmale):

You have correctly used " ^ " to denote exponentiation, but your "x-1" is subject to misinterpretation. I assume (but do not know) that you meant g(x)=3(2)^(x-1), or \[g(x)=3(2)^{x-1}\]

OpenStudy (mathmale):

Clarify this before we move on.

OpenStudy (snickers1998):

Its supposed to be the second choice

OpenStudy (mathmale):

Does\[g(x)=3(2)^{x-1}\] seem to have a maximum value? a minimum value? These questions are relevant to "range."

OpenStudy (mathmale):

Are you able to graph y=2^x? That's an exponential function.

OpenStudy (mathmale):

If unsure, you might want to Google "exponential function." You'll most likely see sample graphs right away.

OpenStudy (mathmale):

Snickers: Where are you?

OpenStudy (snickers1998):

the question just wants to know if all real numbers are greater than or equal to 5 or all real numbers are greater than -1 or all real numbers are greater than or equal to -1

OpenStudy (snickers1998):

I was working on other math problems sorry im going back and forth between screens

OpenStudy (mathmale):

I'd suggest you look up "range of a function." Do that first, please. Then: "Are you able to graph y=2^x? That's an exponential function." What is the range of this sample function?

OpenStudy (mathmale):

I'm signing off for now. Others may be able to continue helping you. I will be back on OpenStudy later today. Good luck!